Suppose that NVIDIA Corporation (NVDA) stock is selling for $150.00. Analysts believe that the growth rate for NVDA will be 20% next year, 30% for the following 4 years, 10% for the following two years, and thereafter the growth rate will be 6% indefinitely. NVDA will pay a cash dividend of $.65 per share next year. Thereafter the dividend will grow by the same rate as the company. Stockholders require a return of 16 percent on NVIDIA’s stock.

Required:

  1. Based on the above assumptions, determine the price of NVIDIA’s common stock.
  2. Explain whether an investor should buy the stock.
Answer #1
A B C D E F G H I J K L M N O P Q
2
3 a)
4 Price of the stock will be the present value of the future dividends.
5 Dividend next year $0.65
6 Required rate of return 16%
7
8 Year 0 1 2 3 4 5 6 7 8 9
9 Dividend growth rate 20% 30% 30% 30% 30% 10% 10% 6% 6% 6%
10 Dividend $0.65 $0.85 $1.10 $1.43 $1.86 $2.04 $2.25 $2.38 $2.52 =J10*(1+K9)
11 Terminal value = DIV8/(rs-gL) $23.81 =L10/(D6-L9)
12 Present value of dividends $0.56 $0.63 $0.70 $0.79 $0.88 $0.84 $9.22 =(K10+K11)/((1+$D$6)^K8)
13 Price of share at Year 0 $13.62 =SUM(E12:K12)
14
15 Thus the price of stock is $13.62
16
17 b)
18
19 Market price of the stock is $150 whereas intrinsic value of stock is $13.60,
20 therefore the share is overpriced.
21 Hence the share should not be bought.
22